OpenStudy (anonymous):

prove by mathematical induction that for n even n^2+ 2n is a multiple of 8

OpenStudy (anonymous):

I did it till this step (k+1) (k+3), what to do after this?

OpenStudy (anonymous):

P(2)=8 == 8*1 ==> P(2) is true Let P(n) be true ==> n^2+2n=8k P(n+2)=(n+2)^2+2*(n+2) = n^2+2n+4n+4+4 =8k+4n+8 =8(k+1)+4n n is even ==>n= 2x =8(k+1)+4*2x =8*(k+1+x) ==>P(n+2) is a multiple of 8 Hence by induction P(n)= n^2+2n is a multiple of 8 for n even

OpenStudy (anonymous):

Let n=2x then, n^2+2n =4x^2 + 4x = 4*x*(x+1) Now, product of any two consecutive integer is always even So, 4*x*(x+1) is always divisible by 8... Thus, n^2+2n is a multiple of 8 if n is even.